# Evolutionary Prediction Games

### v2

@@ D.8 Hard-SVM With Finite Data
Proof in Section D.8.2. The construction is illustrated in Figure 3 (Right), and described formally in Section D.8.1. Informally, the key property we leverage here is asymmetric class variance. In each group, one class has high variance, and the other class has low variance. When the dataset is of size $n$, this creates bias against the high-variance class, and thus convergence towards the optimal classifier at rate $O(n^{-1})$. When the two groups coexist, this bias is balanced, and the rate of convergence is exponential—leading to coexistence that is beneficial, but unstable.
Definition 13 (Hard-SVM; e.g. [45, Section 15.1]).
@@
Remark 1 (One-dimensional Hard-SVM).
@@ D.8.1 Construction
$X_{A}$ $\sim 0.5\left(\mathrm{Uniform}\left([0,1]\right)+\varepsilon\right)+0.5\delta(x=-\varepsilon)$
@@
$Y|X$ =$=\begin{cases}1&X\geq 0\\ -1&X<0\end{cases}$
$D_{A}$ is the distribution of tuples $(X_{A},Y|X_{A})$, and $D_{B}$ is defined correspondingly as $D_{B}=(X_{B},Y|X_{B})$. The distributions are illustrated in Figure 3 (Right).
@@ D.8.2 Mutualistic Coexistence
By eq. 16 we obtain the expectation of $x_{\mathrm{min},n}$:
@@
eq. 15 gives a lower bound on the expectation:
@@
For case (1), the Hard-SVM algorithm trains on data sampled from $D_{A}$. Denote the size of the dataset by $n$ and the corresponding Hard-SVM classfier by $h_{n}$. We note that $h_{n}$ is a random variable depending on the dataset. For a dataset ${\left\{{(x_{i},y_{i})}\right\}}_{i=1}^{n}\sim D_{A}^{n}$, Since the problem is one-dimensional, the Hard-SVM decision margin is given by eq. 14:
@@
By Proposition 19 and for $n\geq 15$, it holds that $\mathbb{E}{\left[{x_{\mathrm{margin}}}\right]}\geq\frac{1}{n}+\varepsilon$, and therefore:
@@ D.9 Stabilizing Coexistence
Proof of Proposition 1.
Let $F({\bm{p}})$ be an evolutionary prediction game induced by an optimal learning algorithm $\mathcal{A}({\bm{p}})$, and let ${\bm{p}}^{*}$ be an unstable coexistence equilibrium with full support (${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}=K$). The game induced by $\mathcal{A}^{\prime}({\bm{p}})$ in the neighborhood of ${\bm{p}}^{*}$ is $F^{\prime}({\bm{p}})=F(2{\bm{p}}^{*}-{\bm{p}})$. By Lemma 1, the game $F({\bm{p}})$ is a potential game, and therefore admits a potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$. Consider the potential function $f^{\prime}({\bm{p}})=-f(2{\bm{p}}^{*}-{\bm{p}})$. By the chain rule, it holds that:
@@
And therefore $f^{\prime}({\bm{p}})$ is a potential function for the game $F^{\prime}({\bm{p}})$ in the neighborhood of ${\bm{p}}^{*}$. The equilibrium ${\bm{p}}^{*}$ is unstable, and therefore by [43, Theorem 8.2.1] it is not a maximizer of $f({\bm{p}})$. Moreover, by Proposition 3 the potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex, and as ${\bm{p}}^{*}$ is assumed to have full support, it is also a global minimizer of $f({\bm{p}})$. From this we conclude that ${\bm{p}}^{*}$ is a global maximizer of $f^{\prime}({\bm{p}})$, and therefore the equilibrium ${\bm{p}}^{*}$ is stable under the evolutionary game induced by $\mathcal{A}^{\prime}$. ∎
Appendix E Experiment Details
@@
We implement our analysis in Python. Our synthetic-data experiments rely on scikit-learn for learning algorithm implementations [38], and our real-data experiments rely on PyTorch [37] and ffcv [29]. We use the matplotlib package for plotting, and the mpltern package for ternary plots. Code is available at: https://github.com/edensaig/evolutionary-prediction-games.
@@
Synthetic data simulations were run on a single Macbook Pro laptop, with 16GB of RAM, M2 processor, and no GPU. Experiments involving neural networks (Section 6) were run on a dedicated server with an AMD EPYC 7502 CPU, 503GB of RAM, and an Nvidia RTX A4000 GPU.
@@
A single run of the complete synthetic data pipeline takes roughly 20 minutes on a laptop. A single repetition of each real-data experiment (i.e. sampling $h\sim\mathcal{A}({\bm{p}})$ and computing marginal accuracies) takes roughly 10 minutes.
@@
For CIFAR-10, we use the Resnet-9 architecture and training code provided in the CIFAR-10 example code in the ffcv Github repository (libffcv/ffcv, commit 7885f40), with modifications to control the probability of horizontal flips in training and testing. For MNIST, we use the convolutional neural network provided in the MNIST example code in the PyTorch examples Github repository (pytorch/examples, commit 37a1866). The MNIST network has two convolutional layers, and two fully-connected layers, with dropout and max pooling. Training is performed for 200 epochs using SGD with learning rate $0.01$ and momentum $0.5$.
@@
where $\bar{F}({\bm{p}})=\sum_{k}p_{k}F_{k}({\bm{p}})$ is the average fitness across the population. In our simulations (e.g., Figures 2, 4), we use the discrete-time replicator equation given by Taylor and Jonker [48, eq. (3)]:
@@
For the CIFAR-10 replicator simulation (Section 6.2), we discretize the two-group simplex into the grid ${\left\{{(1,0),(0.9,0.1),\dots,(0,1)}\right\}}$, precompute fitness values $F_{k}({\bm{p}})$ at each grid point, and use linear interpolation to determine fitness at intermediate states.
@@
In both experiments, variability is due to the random split within the train and test sets, and due to random noise when it is added. For numerical results, we report normal mean confidence intervals at $99\%$ confidence level.

### v3

@@ D.8 Hard-SVM With Finite Data
Proof in Section˜D.8.2. The construction is illustrated in Figure˜3 (Right), and described formally in Section˜D.8.1. Informally, the key property we leverage here is asymmetric class variance. In each group, one class has high variance, and the other class has low variance. When the dataset is of size $n$, this creates bias against the high-variance class, and thus convergence towards the optimal classifier at rate $O(n^{-1})$. When the two groups coexist, this bias is balanced, and the rate of convergence is exponential—leading to coexistence that is beneficial, but unstable.
Definition D.13 (Hard-SVM; e.g. [54, Section 15.1]).
@@
Remark D.1 (One-dimensional Hard-SVM).
@@ D.8.1 Construction
[…]
@@
$Y|X$ $1X<0\displaystyle$=
$D_{A}$ is the distribution of tuples $(X_{A},Y|X_{A})$, and $D_{B}$ is defined correspondingly as $D_{B}=(X_{B},Y|X_{B})$. The distributions are illustrated in Figure˜3 (Right).
@@ D.8.2 Mutualistic Coexistence
By eq.˜16 we obtain the expectation of $x_{\mathrm{min},n}$:
@@
eq.˜15 gives a lower bound on the expectation:
@@
For case (1), the Hard-SVM algorithm trains on data sampled from $D_{A}$. Denote the size of the dataset by $n$ and the corresponding Hard-SVM classfier by $h_{n}$. We note that $h_{n}$ is a random variable depending on the dataset. For a dataset ${\left\{{(x_{i},y_{i})}\right\}}_{i=1}^{n}\sim D_{A}^{n}$, Since the problem is one-dimensional, the Hard-SVM decision margin is given by eq.˜14:
@@
By Proposition˜D.19 and for $n\geq 15$, it holds that $\mathbb{E}{\left[{x_{\mathrm{margin}}}\right]}\geq\frac{1}{n}+\varepsilon$, and therefore:
@@ D.9 Stabilizing Coexistence
Proof of Proposition˜2.
Let $F({\bm{p}})$ be an evolutionary prediction game induced by an optimal learning algorithm $\mathcal{A}({\bm{p}})$, and let $\bm{p}$$^{*}$ be an unstable coexistence equilibrium with full support (${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}=K$). The game induced by $\mathcal{A}^{\prime}({\bm{p}})$ in the neighborhood of $\bm{p}$$^{*}$ is $F^{\prime}({\bm{p}})=F(2{\bm{p}}^{*}-{\bm{p}})$. By Lemma˜D.1, the game $F({\bm{p}})$ is a potential game, and therefore admits a potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$. Consider the potential function $f^{\prime}({\bm{p}})=-f(2{\bm{p}}^{*}-{\bm{p}})$. By the chain rule, it holds that:
@@
And therefore $f^{\prime}({\bm{p}})$ is a potential function for the game $F^{\prime}({\bm{p}})$ in the neighborhood of $\bm{p}$$^{*}$. The equilibrium $\bm{p}$$^{*}$ is unstable, and therefore by [52, Theorem 8.2.1] it is not a maximizer of $f({\bm{p}})$. Moreover, by Proposition˜D.3 the potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex, and as $\bm{p}$$^{*}$ is assumed to have full support, it is also a global minimizer of $f({\bm{p}})$. From this we conclude that $\bm{p}$$^{*}$ is a global maximizer of $f^{\prime}({\bm{p}})$, and therefore the equilibrium $\bm{p}$$^{*}$ is stable under the evolutionary game induced by $\mathcal{A}^{\prime}$. ∎
Appendix E Implementation Details
@@
We implement our simulations and analysis in Python. Our synthetic-data experiments rely on scikit-learn [47] for learning algorithm implementations, our CIFAR-10 and MNIST experiments rely on PyTorch [46] and ffcv [38], and our ACSIncome experiment relies on scikit-learn and XGBoost [10]. We use matplotlib [29] for plotting, and mpltern [30] for ternary plots. Code is available at: https://github.com/edensaig/evolutionary-prediction-games.
@@
Synthetic data simulations were run on a single Macbook Pro laptop, with 16GB of RAM, M2 processor, and no GPU. Experiments involving neural networks (Section˜6) were run on a dedicated server with an AMD EPYC 7502 CPU, 503GB of RAM, and an Nvidia RTX A4000 GPU.
@@
A single run of the complete synthetic data pipeline takes roughly 30 minutes on a laptop. A single repetition of each real-data experiment (i.e. sampling $h\sim\mathcal{A}({\bm{p}})$ and computing marginal accuracies) takes roughly 10 minutes.
@@
For CIFAR-10, we use the Resnet-9 architecture and training code provided in the CIFAR-10 example code in the ffcv Github repository (libffcv/ffcv, commit 7885f40), with modifications to control the probability of horizontal flips in training and testing. For MNIST, we use the convolutional neural network provided in the MNIST example code in the PyTorch examples Github repository (pytorch/examples, commit 37a1866). The MNIST network has two convolutional layers, and two fully-connected layers, with dropout and max pooling. Training is performed for 200 epochs using SGD with learning rate $0.01$ and momentum $0.5$. For the ACSIncome experiments, we use scikit-learn and XGBoost classifiers with default regularization parameters (LinearSVC, LogisticRegression, XGBClassifier).
@@
where $\bar{F}({\bm{p}})=\sum_{k}p_{k}F_{k}({\bm{p}})$ is the average fitness across the population. In our simulations (e.g., Figures 2, 4), we use the discrete-time replicator equation given by Taylor and Jonker [57, eq. (3)]:
@@
For the CIFAR-10 replicator simulation (Section˜6.1), we discretize the two-group simplex into the grid $\left\{{(1,0),(0.9,0.1),\dots,(0,1)}\right\}$}, precompute fitness values $F_{k}({\bm{p}})$ at each grid point, and use linear interpolation to determine fitness at intermediate states.
@@
In both experiments, variability is due to the random split within the train and test sets, and due to random noise when it is added. For numerical results, we report mean confidence intervals at $99\%$ confidence level.
Appendix F Additional Empirical Evaluation
F.1 Effect of Sampling Noise
Figure 12: Effect of sampling noise in the CIFAR-10 setting (Section˜F.1). (Left) Gaussian process fit to the raw observations. Error bars indicate standard deviation of data, and shaded areas indicate standard deviation of the Gaussian process. (Center) Trajectory samples of the induced stochastic replicator dynamics, for different simulated noise levels $\alpha$. (Right) Distribution of long-term outcomes as a function of initial condition.
The prediction games defined in Section˜3 associate evolutionary fitness with accuracy in expectation over the training set (eq.˜2, a reasonable assumption where sampling noise is small. Here we test the robustness of this assumption by quantifying the effect of sampling noise on long-term outcomes.
Method.
We assess the effect of sampling noise on long term outcome by fitting a Gaussian process to the outcomes of different sampled training sets, and analyzing the resulting stochastic dynamics for varying initial conditions. We use the raw data from the CIFAR-10 experiment introduced in Section˜6.1 (group accuracies for each training set sample). We fit a Gaussian process (GP) with an RBF kernel with length scale $0.1$, using the skicit–learn implementation with alpha regularization parameter $4$. To generate stochastic replicator dynamics, we replace the deterministic fitness terms in eq.˜17 with samples from the Gaussian process. We simulate different noise levels by multiplying the Gaussian process covariance by a constant $\eta\geq 0$, such that $\eta=0$ coincides with the deterministic dynamics, $\eta=1$ coincides with the Gaussian process fit to the original data, and $\eta=5$ represents a simulated five-fold increase in compared to the observed sampling noise. For each $\eta$, we simulate stochastic dynamics for a range of initial states $p^{0}_{B}\in[0.35,0.65]$, and measure the distribution of long term outcomes (dominance of group $A$ vs. group $B$) over repeated realizations.
Results.
Gaussian process fit is illustrated in Figure˜12 (Left), stochastic trajectories are presented in Figure˜12 (Center), and distributions of long-term outcomes are presented in Figure˜13 (Right). When there is no sampling noise (i.e., $\eta=0$), group $B$ dominates iff $p^{0}_{B}>0.5$ as expected. With the natural variance of the task (i.e., using the inferred Gaussian process covariance and $\eta=1$), results are relatively robust, and group $B$ eventually dominates with high probability for all $p^{0}_{B}>0.5$. When variance is excessively increased (i.e., $\eta=5$), outcomes also become noisy, and dynamics can converge to group $A$ with some probability even for $p^{0}_{B}\approx 0.55$.
F.2 Time to Convergence
For unstable equilibria, a key question is how fast the system transitions to a stable equilibrium state, and what affects this rate. Here we measure the time to reach (approximate) domination by one group as a function of the initial state.
Method.
We extend the analysis of the CIFAR-10 experiment, described in Section˜6.1. To obtain a smooth population game, we use the mean curves of the Gaussian process described in Section˜F.1. For each initial condition, we simulate the replicator dynamics, and record the number of steps until one group is approximately dominant, formally $\min{\bm{p}}^{t}\leq 0.01$.
Results.
Figure˜13 (Center) presents the measured convergence times. Results show that the time to convergence is roughly linear in $p^{0}_{B}$ for $p^{0}_{B}\in[0.1,0.4]\cup[0.6,0.9]$, but tends to infinity as the initial state approaches the uniform distribution ($p^{0}_{B}\to 0.5)$. The implication is that for initially balanced populations, natural selection forces are relatively weak, and and higher-order effects (such as sampling noise, or other exogenous forces) can dominate the dynamics.
Figure 13: Empirical analysis of deterministic sensitivity aspects (Appendices F.2, F.3). (Left) Evolutionary prediction game induced by CIFAR-10 with horizontal flips and Resnet-9. Curves are smoothed by fitting a Gaussian process, as described in Section˜F.1. (Center) Time to convergence for different initial conditions. Curve is relatively linear far from the unstable equilibrium, but approach infinity as the initial condition approaches the unstable equilibrium. (Right) Sensitivity to equilibrium estimation errors in steering, showing a linear relation.
F.3 Sensitivity Analysis for Stabilization
Our proposed stabilization mechanism (Section˜5.2) assumes access to the true equilibrium $\bm{p}$$^{*}$, which may not be known in practice. Although our experiments in Section˜6.1 use an estimated $\bm{p}$$^{*}$ and results comply with our theoretical findings, they do reveal how robust outcomes are to misspecification of $\bm{p}$$^{*}$. For this, we present a sensitivity analysis quantifying the deviation of the reached state from the estimated equilibrium $\bm{p}$$^{*}$.
Method.
We simulate the stabilization method presented in Section˜5.2 for varying population states around the uniform population state equilibrium, and measure the resulting long-term outcomes of the dynamics.
Results.
Results are presented in Figure˜13 (Right), and show a linear relation between the estimated equilibrium and the eventual outcome of the stabilized dynamics. An additional observation is that misspecification affects only the population composition, and not welfare. This is because the algorithm ensures that the returned classifier acts “as if” the system is under equilibrium, in which accuracy for both groups is the same.
F.4 Three-Group Dynamics for Different Learning Algorithms
To further establish the results of Section˜6.3, we compare three-group dynamics induced by different algorithms in the same setting.
Method.
We follow the same procedure presented in Section˜6.3, and compute the dynamics for a linear SVM (from scikit-learn), logistic regression (scikit-learn), and an XGBoost classifier, all with default regularization parameters. Replicator dynamics are simulated using Equation˜17, and basins of attraction are computed by simulating the dynamics until convergence. The initial condition for all algorithms is ${\bm{p}}^{0}=\left(p_{\mathrm{(TX)}},p_{\mathrm{(NY)}},p_{\mathrm{(CA)}}\right)=(0.375,0.5,0.125)$.
Results.
Results are presented in Figure˜14. For identical initial conditions, each learning algorithm leads to a qualitatively different outcome: dynamics induced by the linear SVM converge to dominance of the Texas (TX) user group, dynamics induced by logistic regression converge towards a coexistence saddle point, and XGBoost dynamics converge towards dominance of the California (CA) user group. In particular, for the logistic regression system, note that the population state evolves along the attraction basin boundary (separatrix). By the definition of the replicator dynamics, states along the attraction basin boundary satisfy $\mathrm{acc}_{\mathrm{(CA)}}(h_{\bm{p}})=\mathrm{acc}_{\mathrm{(TX)}}(h_{\bm{p}})$, and therefore exhibit a two-phase phenomenon similar to the one presented in Section˜6.3 (an early balance followed by competition between the remaining groups). In contrast, all interior points in the XGBoost dynamics evolove toward CA domination, precluding the existence of the two-phase phenomenon in this setting.
Figure 14: Comparing three-group replicator dynamics across different learning algorithms for the ACSIncome prediction task (Section˜F.4). Background colors represent basins of attraction, and red lines represent trajectories starting from the same initial condition.